How do i solve by elimination?
-x+2y=-10
5x+4y=8

Answer by nerdybill(7384)

-x+2y=-10
5x+4y=8
.
Multiply top equation by -2 and combine with second:
2x-4y = 20
5x+4y = 8
--------------
7x = 28
x = 4
.
now, substitute the value found above into either of the two equations to find y:
5x+4y=8
5(4)+4y=8
20+4y=8
4y=-12
y = -3
.
answer: x=4 and y=-3